# Persistence Paradox in Dynamic Science

### v1

@@ 4 Dataset
We did not use these authors’ papers back to the earliest stages of machine learning in the 1980s. Given the generational turnover of scientists and the evolutionary nature of scientific progress, their papers between 2003 and 2012 are more representative of the pre-deep learning scientific community and the immediate reaction of the community to a paradigm shift. These scientists were relatively senior when they encountered the deep learning boom in 2013. We recorded their academic age—calculated as the difference between 2013 and the year of their first publication in the OpenAlex database—to represent their experience at the time of this paradigm shift (Figure 1). On average, they had been active in the field for over a decade, having published in the most prestigious venues during the formative years of 2003–2012.
@@ 5.2.1 Will more adaptive scientists be more successful?
  • Research Persistence (0-100): To quantify a scientist’s research persistence from year $T$ to $T$ + 1, we first gather all ICML/NeurIPS papers authored by the scientist up to and including year $T$. Next, we collect their ICML/NeurIPS papers from year $T$ + 1. To assess persistence, we compute the bootstrapped text similarity between these two corpora. Since the sizes of the two corpora (before $T$ and $T$ + 1) may differ significantly—introducing potential bias (text similarity is typically lower when comparing a short text to a longer one)—we address this by matching their sizes. Specifically, we retain the smaller corpus and randomly sample a subset of texts from the larger corpus to match the size of the smaller corpus. We then calculate the text similarity between the sampled subset and the smaller corpus. This process is repeated 1,000 times, and the average similarity across iterations is used as the similarity score from year $T$ to $T$ + 1. A higher bootstrapped similarity score indicates greater persistence and less adaptation.
@@
  • Collaboration Familiarity: For each scientist, we track their unique collaborators in a given year and calculate the number of prior ICML/NeurIPS papers they coauthored with these collaborators (dating back to the inception of these conferences: ICML 1980 and NeurIPS 1987). The scientist’s collaboration age for that year is defined as the average number of prior collaborations with their current collaborators.
@@
Descriptive statistics for the variables presented can be found in Figure 3. It reveals a long-tail distribution in productivity and collaboration networks. Interestingly, there is generally low persistence observed among authors in the field of computer science. The relationship between variables of persistence, prior status, age, and collaboration network features is shown in Appendix D.
Figure 3: Descriptive statistics for variables used in the regression.
@@
Collaborator familiarity, count, and status exhibit mixed effects on productivity and impact. Having more collaborators enhances productivity and impact. Collaborating with high-status partners appears to significantly enhance influence. However, this advantage comes with a tradeoff, as such collaborations tend to negatively affect productivity, potentially due to the time and resources required to maintain these high-status relationships and high-status authors’ high standard of publishing. Teams with older collaborations exhibit lower productivity and impact levels. Old teams are less successful in adapting to the new paradigm. We further examined the ”age” [^6] This age is the age of a team. In the paper we also have the age of a scientist, defined as the years since their first publication. of these scientific teams as shown in Figure 4, defined as the average number of ICML/NeurIPS papers the focal author had previously collaborated on with team members that year (right), and the proportion of new collaborations, defined as the percentage of team members who had never co-authored an ICML/NeurIPS paper with the focal author before (left). We tracked their collaborations back to the initial year of these conferences. We found that right after the shock, published teams tended to be more experienced — presumably those at the center of deep learning prior to the shock - and they continued to adapt through collaborating with newcomers in the years following 2013. This is evidenced by a decreasing average team age and an increasing share of new collaborators, both of which contributed to the continued success. Notably, these new collaborations tended to persist over time, as teams gradually became old again in subsequent years. These effects are robust within individual conferences, see Appendix B.
@@ 5.3 Summary of Results
Interestingly, our results suggest that the motivations and capacities for adaptation vary across different strata of the scientific community. Low-status scientists, constrained entrenched positions, may feel a heightened urgency to explore new directions during periods of disruption. This allows them to capitalize on emerging opportunities and challenge the existing power dynamics. This interplay of disruption, status, and adaptability underscores the complexity of scientific evolution, where both the persistence of existing paradigms and the emergence of novel directions coexist and compete.
@@ Appendix B Teams adapt through fresh collaborations within the individual conference
As shown in Figures 9 and 10 the patterns reported in the main paper (Figure 4) are robust within individual conferences.
@@ Appendix C The impact of ICLR
The second possible channel is that some scientists shifted their focus to publishing at ICLR so they disappeared from ICML/NeurIPS. However, only 1.21% (65) of the focal group of scientists submitted at least one paper to ICLR 2013, yet they accounted for 65.7% of submissions and 77.3% of accepted papers at the conference. During the first five years of ICLR (2013–2017) - a period marked by heightened dynamism as scientists formed new collaborations to adapt (see Figures 4, 9, and 10) - papers published in ICLR accounted for only 6.8% of this cohort’s total publications. Although ICLR later grew in prominence, it is important to recognize that such shifts in impact take time to materialize.
@@
Appendix D Relationships between previous status, collaboration networks, and age
@@
Appendix E Regressions of age and productivity/impact
@@
Impact has a negative relation with age but the association is very weak (coefficient $=-0.0002$ (in a 0-100 scale), p = 0.048). Productivity actually has a strong positive relation with age (coefficient = 0.007, p = 0.000).

### v2

@@ 4 Dataset
We did not use these authors’ papers back to the earliest stages of machine learning in the 1980s. Given the generational turnover of scientists and the evolutionary nature of scientific progress, their papers between 2003 and 2012 are more representative of the pre-deep learning scientific community and the immediate reaction of the community to a paradigm shift. These scientists were relatively senior when they encountered the deep learning boom in 2013. We recorded their academic age—calculated as the difference between 2013 and the year of their first publication in the OpenAlex database—to represent their experience at the time of this paradigm shift (Figure 1)  [^5] This age is the academic seniority of a scientist. Note that in the following sections, we also have the ”team age” (i.e., collaboration familiarity), defined as the average number of previous collaborations on ICML/NeurIPS papers between the focal scientists and their coauthors, see Section 5.2.. On average, they had been active in the field for over a decade, having published in the most prestigious venues during the formative years of 2003–2012.
@@ 5.2.1 Will more adaptive scientists be more successful?
  • Research Persistence (0-100): To quantify a scientist’s research persistence from year $T$ to $T$ + 1, we first gather all ICML/NeurIPS papers authored by the scientist up to and including year $T$. Next, we collect their ICML/NeurIPS papers from year $T$ + 1. To assess persistence, we compute the bootstrapped text similarity between these two corpora. Since the sizes of the two corpora (before $T$ and $T$ + 1) may differ significantly—introducing potential bias (text similarity is typically lower when comparing a short text to a longer one)—we address this by matching their sizes. Specifically, we retain the smaller corpus and randomly sample a subset of texts from the larger corpus to match the size of the smaller corpus. We then calculate the text similarity between the sampled subset and the smaller corpus. This process is repeated 1,000 times, and the average similarity across iterations is used as the similarity score from year $T$ to $T$ + 1. A higher bootstrapped similarity score indicates greater persistence and less adaptation. Interestingly, there is generally low persistence (41 $\pm$ 1 on a 0-100 scale) observed among authors in the field of computer science.
@@
  • Collaboration Familiarity: For each scientist, we track their unique collaborators in a given year and calculate the number of prior ICML/NeurIPS papers they coauthored with these collaborators (dating back to the inception of these conferences: ICML 1980 and NeurIPS 1987). The scientist’s collaboration familiarity with their coauthors for that year (i.e., team age) is defined as the average number of prior collaborations with their current collaborators.
@@
The relationship between variables of persistence, prior status, academic age, and collaboration network features is shown in Appendix D.
@@
Collaborator familiarity, count, and status exhibit mixed effects on productivity and impact. Having more collaborators enhances productivity and impact. Collaborating with high-status partners appears to significantly enhance influence. However, this advantage comes with a tradeoff, as such collaborations tend to negatively affect productivity, potentially due to the time and resources required to maintain these high-status relationships and high-status authors’ high standard of publishing. Teams with older collaborations exhibit lower productivity and impact levels. Old teams are less successful in adapting to the new paradigm. We further examined the ”age” of these scientific teams as shown in Figure 3, defined as the average number of ICML/NeurIPS papers the focal author had previously collaborated on with team members that year (right), and the proportion of new collaborations, defined as the percentage of team members who had never co-authored an ICML/NeurIPS paper with the focal author before (left). We tracked their collaborations back to the initial year of these conferences. We found that right after the shock, published teams tended to be more experienced — presumably those at the center of deep learning prior to the shock - and they continued to adapt through collaborating with newcomers in the years following 2013. This is evidenced by a decreasing average team age and an increasing share of new collaborators, both of which contributed to the continued success. Notably, these new collaborations tended to persist over time, as teams gradually became old again in subsequent years. These effects are robust within individual conferences, see Appendix B.
@@ 5.3 Summary of Results
Interestingly, our results suggest that the motivations and capacities for adaptation and persistence vary across different strata of the scientific community. Low-status scientists, constrained entrenched positions, may feel a heightened urgency to explore new directions during periods of disruption. This allows them to capitalize on emerging opportunities and challenge the existing power dynamics. This interplay of disruption, status, and adaptability underscores the complexity of scientific evolution, where both the persistence of existing paradigms and the emergence of novel directions coexist and compete.
@@ Appendix B Teams adapt through fresh collaborations within the individual conference
As shown in Figures 8 and 9 the patterns reported in the main paper (Figure 3) are robust within individual conferences.
@@ Appendix C The impact of ICLR
The second possible channel is that some scientists shifted their focus to publishing at ICLR so they disappeared from ICML/NeurIPS. However, only 1.21% (65) of the focal group of scientists submitted at least one paper to ICLR 2013, yet they accounted for 65.7% of submissions and 77.3% of accepted papers at the conference. During the first five years of ICLR (2013–2017) - a period marked by heightened dynamism as scientists formed new collaborations to adapt (see Figures 3, 8, and 9) - papers published in ICLR accounted for only 6.8% of this cohort’s total publications. Although ICLR later grew in prominence, it is important to recognize that such shifts in impact take time to materialize.
@@
Appendix D Relationships between previous status, collaboration networks, and academic age
@@
Appendix E Regressions of academic age and productivity/impact
@@
Impact has a negative relation with age but the association is very weak (coefficient $=-0.0002$ (in a 0-100 scale), p = 0.048). Productivity actually has a strong positive relation with age (coefficient = 0.007, p = 0.000). Here age is at the scientist level not the team level.